Slope of function if y=20x^3+0.5 when x=3


1
Expert's answer
2020-11-10T07:24:29-0500
"y=20x^3+0.5"


"y^{/}=20\times 3x^2+0=60x^2"


"y^/(3)=60\times 3^2=540"